Claudia has sold her car and received approval from the garage owner to re-lease her downtown reserved parking spot for the next four months so she can make some extra money. The rental fee is $200 per month, and she expects to charge $18 per day. Transportation in a car pool will cost her $6 per day. If there are a maximum of 20 work days per month for re-leasing the spot, determine the following: a. Total cost and revenue relations b. Breakeven quantity per month c. Amount of money she will make (or lose) if the number of re-leased days per month over the four-month period are 18, 12, 17, and 20.


1
Expert's answer
2021-10-27T08:30:32-0400

(a)

Total cost=rental fee+ transportation fee.

In a month:

"=200+(6\times20)=320."

In 4 months:

"=320\times4"

=$1280

Total revenue in a month: "=18\times20=360."

In 4 months:

"360\times4=1440"

=$1440.


(b)

Break even quantity per month:

"=\frac{Fixed Cost}{P-VC}"

"=\frac{200}{(18\times20)-(6\times20)}"

"=0.8"

(c)

Profit for 18 days:

Total cost ="200+(6\times18)=308."

Total revenue"=18\times18=324"

Profit "=324-308=16"


For 12 days:

Total cost"=200+(6\times12)=272"

Total revenue"=18\times12=216"

Here, a loss of $56 is made.


For 17 days:

Total cost "=200+(6\times17)=302"

Total revenue"=18\times17=306."

Profit "=306-302=4"


For 20 days:

Total cost "=200+(6\times20)=320."

Total revenue"=18\times20=360."

Profit "=360-320=40"

=$40.